OpenStudy (anonymous):

f(x)= ^3√x-9 find f^-1(x)

OpenStudy (anonymous):

Is that the inverse of a cube root function?

OpenStudy (anonymous):

idk thats how it was written. simplify answer

OpenStudy (anonymous):

I'm trying to check how you have typed it up. It's not totally clear to me, but assuming that is.... To find the inverse you switch the x and y terms and simplify.

OpenStudy (anonymous):

\[ f(x)=\sqrt[3]{x-9} \]To find \(f^{-1}(x)\) we replace \(f(x)\) with \(x\) and \(x\) with \(f^{-1}(x)\):\[ x=\sqrt[3]{f^{-1}(x)-9} \]Then you must solve for \(f^{-1}(x)\).

OpenStudy (anonymous):

Cube both sides:\[ x^3=f^{-1}(x)-9 \]Add 9 to both sides:\[ f^{-1}(x) = x^3+9 \]

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

What is it?